# Transformers are Minimax Optimal Nonparametric In-Context Learners

### v1

@@ 1 Introduction
The paper is structured as follows. In Section 2, the regression tasks and transformer model are defined in an abstract setting. In Section 3, we present the general framework for estimating the ICL approximation and generalization error. In Section 4, we specialize to the Besov-type and piecewise $\gamma$-smooth class settings and show that transformers can achieve or exceed the minimax optimal rate in context. In Section 5, we derive minimax lower bounds. All proofs are deferred to the appendix; moreover, we provide numerical experiments validating our results in Appendix F.
@@ 5 Minimax Lower Bounds
For this section we assume that the noise (1) is i.i.d. Gaussian, $\xi_{k}\sim\mathcal{N}(0,\sigma^{2})$, instead of bounded; while the exact shape of the noise distribution is not important, having restricted support may convey additional information and affect the minimax rate. We also suppose for simplicity that the support of $\operatorname{\mathcal{P}}_{\beta}$ is included in :=${\mathcal{B}}:=\{\beta\in\operatorname{\mathbb{R}}^{\infty}\mid\left\lvert\beta_{j}\right\rvert\lesssim j^{-2s-1}(\log j)^{-2},\;j\in\operatorname{\mathbb{N}}\}$ and that the aggregated coefficients $\bar{\beta}_{j}$ for $\underaccent{\bar}{N}\leq j\leq\bar{N}$ satisfy $\mathbb{E}_{\beta}[\bar{\beta}_{j}^{2}]\leq\sigma_{\beta}^{2}$ for some $\sigma_{\beta}$ dependent on $N$. The proof of the following statement is given in Appendix E.1.
@@
Finally, Proposition 5.1 is applied to obtain concrete lower bounds for the settings studied in Section 4 throughout Appendices E.2-E.4.
@@ Acknowledgments
JK was partially supported by JST CREST (JPMJCR2015). TS was partially supported by JSPS KAKENHI (20H00576) and JST CREST (JPMJCR2115).
@@ C.1.1 Verification of Assumptions
$\lVert F^{\circ}_{\beta}-F^{\circ}_{\beta,\bar{N}}\rVert_{L^{2}(\operatorname{\mathcal{P}}_{\operatorname{\mathcal{X}}})}^{2}$ $=\int_{\operatorname{\mathcal{X}}}\bigg{(}\sum_{j=\bar{N}+1}^{\infty}\beta_{j}\psi_{j}^{\circ}(x)\bigg{)}^{2}\operatorname{\mathcal{P}}_{\operatorname{\mathcal{X}}}(\mathop{}\!\mathrm{d}x)$
$=\sum_{j,k=\bar{N}+1}^{\infty}\beta_{j}\beta_{k}\mathbb{E}_{x}[\psi^{\circ}_{j}(x)\psi^{\circ}_{k}(x)]$
$\leq\lim_{M\to\infty}C_{2}\lVert\beta_{\bar{N}:M}\rVert^{2}$
$=C_{2}\sum_{k=K}^{\infty}\sum_{\ell\in I_{k}^{d}}\beta_{k,\ell}^{2}$
$\lesssim\sum_{k=K}^{\infty}|I_{k}^{d}|^{1-2/p}\Bigg{(}\sum_{\ell\in I_{k}^{d}}|\beta_{k,\ell}|^{p}\Bigg{)}^{2/p}$
$\lesssim\sum_{k=K}^{\infty}2^{kd}\cdot(2^{kd})^{-2\alpha/d-1}$
@@
$\mathbb{E}_{\beta}[\bar{\beta}_{K,\ell}^{2}]$ $=\sum_{k=0}^{K}\sum_{\ell\in I_{k}^{d}}\gamma_{k,K,\ell,\ell^{\prime}}^{2}\mathbb{E}_{\beta}[\beta_{k,\ell}^{2}]$
$\lesssim\sum_{k=0}^{K}\Bigg{(}\sum_{\ell\in I_{k}^{d}}\gamma_{k,K,\ell,\ell^{\prime}}\Bigg{)}^{2}2^{-k(2\alpha+d)}k^{-2}$
@@
$\lesssim 2^{-Kd}\cdot\sum_{k=0}^{K}2^{-2k\alpha}k^{-2}$
$\asymp N^{-1},$
@@ C.2.1 Definitions and Results
$\lVert F^{\circ}_{\beta}-F^{\circ}_{\beta,\bar{N}}\rVert_{L^{2}(\operatorname{\mathcal{P}}_{\operatorname{\mathcal{X}}})}^{2}$ $\lesssim\sum_{k=K+1}^{\infty}\sum_{\ell\in I_{k}^{d,\alpha}}\beta_{k,\ell}^{2}$
$\lesssim\sum_{k=K+1}^{\infty}|I_{k}^{d,\alpha}|^{1-2/p}\Bigg{(}\sum_{\ell\in I_{k}^{d,\alpha}}|\beta_{k,\ell}|^{p}\Bigg{)}^{2/p}$
$\lesssim\sum_{k=K+1}^{\infty}2^{\lVert k\rVert_{\underline{\alpha}/\alpha}}\cdot 2^{-\lVert k\rVert_{\underline{\alpha}/\alpha}-2k\underline{\alpha}}$
$\asymp 2^{-2K\underline{\alpha}}\asymp N^{-2\widetilde{\alpha}}.$
@@
$\mathbb{E}_{\beta}[\bar{\beta}_{K,\ell}^{2}]$ $=\sum_{k=0}^{K}\sum_{\ell\in I_{k}^{d,\alpha}}\gamma_{k,K,\ell,\ell^{\prime}}^{2}\mathbb{E}_{\beta}[\beta_{k,\ell}^{2}]$
$\lesssim\sum_{k=0}^{K}\Bigg{(}\sum_{\ell\in I_{k}^{d,\alpha}}\gamma_{k,K,\ell,\ell^{\prime}}\Bigg{)}^{2}2^{-k\underline{\alpha}(2+1/\widetilde{\alpha})}k^{-2}$
@@
$\lesssim 2^{-\lVert K\rVert_{\underline{\alpha}/\alpha}}\cdot\sum_{k=0}^{K}2^{-2k\underline{\alpha}}k^{-2}$
$\asymp N^{-1},$
@@ D.1 Definitions and Results
Appendix E Proofs of Minimax Lower Bounds
E.1 Proof of Proposition 5.1
@@ Appendix E Proofs of Minimax Lower Bounds
Applying Proposition E.1 together with (7) concludes the proof. ∎
E.2 Lower Bound in Besov Space
@@
E.3 Lower Bound with Coarser Basis
@@
E.4 Lower Bound in Piecewise $\gamma$-smooth Class
@@
Appendix F Numerical Experiments
In this section, we connect our theoretical contributions to practical transformers by conducting numerical experiments verifying our results as well as justifying the simplified model setup and the assumption of empirical risk minimization. We implement and compare the following 3 toy models: (a) the simplified architecture studied in our paper; (b) the same model with linear attention replaced by softmax; and (c) a full transformer with 2 stacked encoder layers. The number of feedforward layers, widths of hidden layers, learning rate, etc. are set to equal for a fair comparison, see Figure 1 for details.
Figure 2 shows training (solid) and test loss curves (dashed) during pretraining. All 3 architectures exhibit similar behavior and converge to near zero training loss, justifying the use of our simplified model and also supporting the assumption that the empirical risk is minimized. Moreover, Figure 3 shows the converged losses over a wide range of $N,n,T$ values. We verify that increasing $N,n$ leads to decreasing train and test error, corresponding to the approximation error of Theorem 3.1. We also observe that increasing $T$ tends to improve the pretraining generalization gap up to a threshold, confirming our theoretical analysis of task diversity. Again, this behavior is consistent across the 3 architectures.
We note that in the overparametrized regime when the number of total parameters $\gtrsim nT$, the trained model is likely not the empirical risk minimizer, which may also contribute to the large error for large $N$ or small $n,T$.

### v2

@@ 1 Introduction
The paper is structured as follows. In Section 2, the regression tasks and transformer model are defined in an abstract setting. In Section 3, we present the general framework for estimating the ICL approximation and generalization error. In Section 4, we specialize to the Besov-type and piecewise $\gamma$-smooth class settings and show that transformers can achieve or exceed the minimax optimal rate in context. In Section 5, we derive minimax lower bounds. All proofs are deferred to the appendix; moreover, we provide numerical experiments validating our results in Appendix E.
@@ 5 Minimax Lower Bounds
For this section we assume that the noise (1) is i.i.d. Gaussian, $\xi_{k}\sim\mathcal{N}(0,\sigma^{2})$, instead of bounded; while the exact shape of the noise distribution is not important, having restricted support may convey additional information and affect the minimax rate. We also suppose for simplicity that the support of $\operatorname{\mathcal{P}}_{\beta}$ is included in :=${\mathcal{B}}:=\{\beta\in\operatorname{\mathbb{R}}^{\infty}\mid\left\lvert\beta_{j}\right\rvert\lesssim j^{-2s-1}(\log j)^{-2},\;j\in\operatorname{\mathbb{N}}\}$ and that the aggregated coefficients $\bar{\beta}_{j}$ for $\underaccent{\bar}{N}\leq j\leq\bar{N}$ satisfy $\mathbb{E}_{\beta}[\bar{\beta}_{j}^{2}]\leq\sigma_{\beta}^{2}$ for some $\sigma_{\beta}$ dependent on $N$. The proof of the following statement is given in Appendix F.1.
@@
Finally, Proposition 5.1 is applied to obtain concrete lower bounds for the settings studied in Section 4 throughout Appendices F.2-F.4.
@@ Acknowledgments
JK was partially supported by JST CREST (JPMJCR2015). TS was partially supported by JSPS KAKENHI (20H00576) and JST CREST (JPMJCR2115). We would like to express our gratitude to Masaaki Imaizumi for valuable and insightful discussions on the topic in relation to his concurrent work in progress (imaizumi2024).
@@ C.1.1 Verification of Assumptions
$\lVert F^{\circ}_{\beta}-F^{\circ}_{\beta,\bar{N}}\rVert_{L^{2}(\operatorname{\mathcal{P}}_{\operatorname{\mathcal{X}}})}^{2}$ $=\int_{\operatorname{\mathcal{X}}}\bigg{(}\sum_{j=\bar{N}+1}^{\infty}\beta_{j}\psi_{j}^{\circ}(x)\bigg{)}^{2}\operatorname{\mathcal{P}}_{\operatorname{\mathcal{X}}}(\mathop{}\!\mathrm{d}x)=\sum_{j,k=\bar{N}+1}^{\infty}\beta_{j}\beta_{k}\mathbb{E}_{x}[\psi^{\circ}_{j}(x)\psi^{\circ}_{k}(x)]$
$\leq\lim_{M\to\infty}C_{2}\lVert\beta_{\bar{N}:M}\rVert^{2}=C_{2}\sum_{k=K}^{\infty}\sum_{\ell\in I_{k}^{d}}\beta_{k,\ell}^{2}$
$\lesssim\sum_{k=K}^{\infty}|I_{k}^{d}|^{1-2/p}\Bigg{(}\sum_{\ell\in I_{k}^{d}}|\beta_{k,\ell}|^{p}\Bigg{)}^{2/p}\lesssim\sum_{k=K}^{\infty}2^{kd}\cdot(2^{kd})^{-2\alpha/d-1}$
@@
$\mathbb{E}_{\beta}[\bar{\beta}_{K,\ell}^{2}]$ $=\sum_{k=0}^{K}\sum_{\ell\in I_{k}^{d}}\gamma_{k,K,\ell,\ell^{\prime}}^{2}\mathbb{E}_{\beta}[\beta_{k,\ell}^{2}]\lesssim\sum_{k=0}^{K}\Bigg{(}\sum_{\ell\in I_{k}^{d}}\gamma_{k,K,\ell,\ell^{\prime}}\Bigg{)}^{2}2^{-k(2\alpha+d)}k^{-2}$
@@
$\lesssim 2^{-Kd}\cdot\sum_{k=0}^{K}2^{-2k\alpha}k^{-2}\asymp N^{-1},$
@@ C.2.1 Definitions and Results
$\lVert F^{\circ}_{\beta}-F^{\circ}_{\beta,\bar{N}}\rVert_{L^{2}(\operatorname{\mathcal{P}}_{\operatorname{\mathcal{X}}})}^{2}$ $\lesssim\sum_{k=K+1}^{\infty}\sum_{\ell\in I_{k}^{d,\alpha}}\beta_{k,\ell}^{2}\lesssim\sum_{k=K+1}^{\infty}|I_{k}^{d,\alpha}|^{1-2/p}\Bigg{(}\sum_{\ell\in I_{k}^{d,\alpha}}|\beta_{k,\ell}|^{p}\Bigg{)}^{2/p}$
$\lesssim\sum_{k=K+1}^{\infty}2^{\lVert k\rVert_{\underline{\alpha}/\alpha}}\cdot 2^{-\lVert k\rVert_{\underline{\alpha}/\alpha}-2k\underline{\alpha}}\asymp 2^{-2K\underline{\alpha}}\asymp N^{-2\widetilde{\alpha}}.$
@@
$\mathbb{E}_{\beta}[\bar{\beta}_{K,\ell}^{2}]$ $=\sum_{k=0}^{K}\sum_{\ell\in I_{k}^{d,\alpha}}\gamma_{k,K,\ell,\ell^{\prime}}^{2}\mathbb{E}_{\beta}[\beta_{k,\ell}^{2}]\lesssim\sum_{k=0}^{K}\Bigg{(}\sum_{\ell\in I_{k}^{d,\alpha}}\gamma_{k,K,\ell,\ell^{\prime}}\Bigg{)}^{2}2^{-k\underline{\alpha}(2+1/\widetilde{\alpha})}k^{-2}$
@@
$\lesssim 2^{-\lVert K\rVert_{\underline{\alpha}/\alpha}}\cdot\sum_{k=0}^{K}2^{-2k\underline{\alpha}}k^{-2}\asymp N^{-1},$
@@ D.1 Definitions and Results
Appendix E Numerical Experiments
In this section, we connect our theoretical contributions to practical transformers by conducting experiments verifying our results as well as justifying the simplified model setup and the empirical risk minimization assumption. We implement and compare the following toy models: (a) the simplified architecture studied in our paper; (b) the same model with linear attention replaced by softmax; and (c) a full transformer with 2 stacked encoder layers. The number of feedforward layers, widths of hidden layers, learning rate, etc. are set to equal for a fair comparison, see Figure 1 for details.
Figure 2 shows training (solid) and test loss curves (dashed) during pretraining. All 3 architectures exhibit similar behavior and converge to near zero training loss, justifying the use of our simplified model and also supporting the assumption that the empirical risk is minimized. Moreover, Figure 3 shows the converged losses over a wide range of $N,n,T$ values. We verify that increasing $N,n$ leads to decreasing train and test error, corresponding to the approximation error of Theorem 3.1. We also observe that increasing $T$ tends to improve the pretraining generalization gap up to a threshold, confirming our theoretical analysis of task diversity. Again, this behavior is consistent across the 3 architectures. We note that in the overparametrized regime when the number of total parameters $\gtrsim nT$, the trained model is likely not the empirical risk minimizer, which may also contribute to the large error for large $N$ or small $n,T$.
Appendix F Proofs of Minimax Lower Bounds
F.1 Proof of Proposition 5.1
@@ Appendix F Proofs of Minimax Lower Bounds
Applying Proposition F.1 together with (7) concludes the proof. ∎
F.2 Lower Bound in Besov Space
@@
F.3 Lower Bound with Coarser Basis
@@
F.4 Lower Bound in Piecewise $\gamma$-smooth Class